# Hardware-efficient Customized Keyword Spotting with Spectral-Temporal Graph Attentive Pooling and a hybrid Loss

### v1

@@ 4 Results and Discussion
Table 1 summarizes FRR of different models at $0.3$ FAs/Hr. In the single word loss configuration, the AAM loss significantly outperforms the CE loss across different encoders. Specifically, AAM improves FRR by $25.7\%$ for ECAPA_TDNN, $31\%$ for Conformer, and $26.1\%$ for LiCoNet. In the hybrid loss configuration featuring the word AAM loss, the inclusion of the reverse speaker loss greatly decreases FRR, particularly for Conformer, resulting in a reduction of $40.9\%$. By incorporating the phoneme loss, we can notice additional enhancements. The efficacy of the hybrid loss underscores the value of using complementary information from both auxiliary losses for KWS modeling.
Table 2: Model size and computation cost of each encoder.
@@
LicoNet $694.1$K $46.5$M
As for the feature aggregator, GAP delivers further substantial improvements compared to ASP across all encoders. In particular, FRR has been decreased by $18.5\%$ for ECAPA_TDNN, $51.1\%$ for Conformer, and $32.2\%$ for LiCoNet. These improvements align with the enhancements observed in the speaker verification task [5] and can be attributed to the increased discriminative capability introduced by the graph pooling strategy. Furhtermore, the SoftTriplet loss effectively captures potential unseen intra-variance within the evaluation data. In the hybrid loss configuration employing graph pooling, the word SoftTriplet loss consistently leads to the best system performance across all models, with a particularly impressive $45.4\%$ reduction in FRR for LiCoNet.
It is noteworthy to see that Conformer consistently maintains superior performance across various loss formulations and pooling strategies. However, LiCoNet achieves comparable performance to Conformer when utilizing the hybrid loss incorporating the word SoftTriplet loss and graph pooling strategy. In Fig. 2, we present DET curves of Conformer using various loss formulations and pooling strategies (left), and those of different encoders while using the hybrid loss featuring SoftTriplet and GAP (right).
Model Efficiency
As shown in Table 2, Conformer boasts the largest model size with considerably more computational demands, despite its superiority in terms of high model capacity. Conversely, LiCoNet strikes a favorable balance between model efficiency and effectiveness, offering performance on par with Conformer while keeping computational costs close to that of ECAPA_TDNN.
@@ 5 Conclusion
In this study, we introduce a hardware-efficient customized KWS system that is centered around the LiCoNet architecture and complemented by a spectral-temporal graph pooling layer and a hybrid loss function. The experimental results showcase the advantages of our framework featuring GAP and the hybrid loss: LiCoNet achieves performance levels similar to those of the computationally intensive Conformer.

### v2

@@ 3.3 Results and Discussion
Single Loss vs. Hybrid Loss In single task learning for word classification, the AAM loss significantly outperforms the CE loss across different encoders. Specifically, AAM improves FRR by $25.7\%$ for ECAPA_TDNN, $31\%$ for Conformer, and $26.1\%$ for LiCoNet. In the hybrid loss configuration featuring the word AAM loss, the inclusion of the reverse speaker loss greatly decreases FRR, particularly for Conformer, resulting in a reduction of $40.9\%$. By incorporating the phoneme loss, we can notice additional enhancements. The efficacy of the hybrid loss underscores the value of using complementary information from both speakers and phonemes for QbyE KWS.
ASP vs. GAP In the hybrid loss setup using word AAM loss, GAP delivers further substantial improvements compared to ASP across all encoders. In particular, FRR has been decreased by $18.5\%$ for ECAPA_TDNN, $51.1\%$ for Conformer, and $32.2\%$ for LiCoNet. These improvements align with the enhancement observed in speaker verification [13] and imply the increased discriminative capability introduced by the graph pooling strategy.
AAM vs. SoftTriplet Further in the hybrid loss configuration employing GAP, the word SoftTriplet loss consistently leads to the best system performance across all models, with a particularly impressive $45.4\%$ reduction in FRR for LiCoNet. This demonstrates the generalizability of SoftTriplet to capture potential unseen intra-variance within the evaluation data.
Encoder Effectiveness Conformer consistently maintains superior performance across various loss formulations and pooling strategies. However, despite the inherent capacity limitations of linear operators in LiCoNet when compared to the attention scheme in Conformer, LiCoNet achieves comparable performance to Conformer ($1.98\%$ vs. $1.63\%$ FRR) in the best multi-task learning setup that employs the word SoftTriplet loss and GAP. This observation demonstrates the effectiveness of the proposed QbyE framework in maximizing the potential of simpler models, making them capable of delivering results on par with their more complex counterparts.
Table 2: Encoder model size and computation on a $2$s audio.
@@
LicoNet $694$K $46.5$M
Model Efficiency To assess memory and computation efficiency, we present the model size and computational cost (FLOPs) in Table 2. It is evident that ECAPA_TDNN exhibits the highest efficiency but has limited performance. On the other hand, Conformer boasts the largest model size with considerably more computational demands, despite its superiority on high model capacity. LiCoNet strikes a favorable balance between model efficiency and effectiveness. It achieves performance on par with Conformer while maintaining computational costs similar to ECAPA_TDNN. Note that the Conformer model size is considerably larger than the other two models. This results from a trade-off between model size and performance. We observe a significant performance degradation when reducing the size of Conformer, whereas only slight performance improvements are observed when increasing the size of ECAPA_TDNN and LiCoNet.
@@ 4 Conclusion
In this study, we introduce a novel QbyE KWS system that employs a spectral-temporal graph pooling layer and multi-task learning. This framework aims to effectively learn speaker-invariant and linguistic-informative embeddings for QbyE KWS tasks. Within this framework, we investigate three distinct network architectures for encoder modeling: LiCoNet, Conformer and ECAPA_TDNN. The experimental results showcase the effectiveness of the proposed QbyE framework in maximizing the potential of simpler models such as LiCoNet, making them capable of delivering results on par with their more complex counterparts.